NH3 is a weak base (Kb = 1.8 × 10-5) and so the salt NH4Cl acts as a weak acid. What is the pH of a solution that is 0.018 M in NH4Cl at 25 C?
Answer #1

The dissociation of ammonium ion in water is as

NH4+ + H2O  --> NH3  + H3O+

The dissociation constant, Ka can be calculated as

Ka*Kb = Kw

Ka = Kw/Kb

= (1.0 x 10-14)/(1.8 x 10-5)

Ka = 5.5 x 10-10

+ H2O -> NH, + H,0* NH, Initial 0.018 M Change -X Equilibrium (0.018-x) +x

The Ka for the above reaction can be calculated as

Ka = ([NH3][H3O+])/[NH4+]

(5.5 x 10-10) = [(x)(x)]/(0.018-x)

x2 + (5.5 x 10-10)x - (9.9 x 10-12) = 0

solving this equation, x = 3.15 x 10-6 M

[H3O+] = x = 3.15 x 10-6 M

Calculate the pH of the solution as follows

pH = -log[H3O+] = -log(3.15 x 10-6)

pH = 5.50 (Ans)